If the average of three numbers is V. If one of the numbers is Z and another is Y, what is the remaining number?
Answer by sdmmadam@yahoo.com(530)

If the average of three numbers is V. If one of the numbers is Z and another is Y, what is the remaining number?
There are three numbers: Z, Y and an unknown R
The average of the three numbers is V
Therefore (Z+Y+R)/3 = V { here Z,Y and V are known)
Multiplying by 3
(Z+Y+R) = 3V
R = (3V-Z-Y)
The remaining number is given by (3V-Z-Y)
Verification: The average of the three numbers = [Z+y+(3V-Z-Y)]/3 = 3V/3 = V which is correct.